# GlycanML: A Multi-Task and Multi-Structure Benchmark for Glycan Machine Learning

### v1

@@ Abstract
Glycans are basic biomolecules and perform essential functions within living organisms. The rapid increase of functional glycan data provides a good opportunity for machine learning solutions to glycan understanding. However, there still lacks a standard machine learning benchmark for glycan function prediction. In this work, we fill this blank by building a comprehensive benchmark for Glycan Machine Learning (GlycanML). The GlycanML benchmark consists of diverse types of tasks including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. Glycans can be represented by both sequences and graphs in GlycanML, which enables us to extensively evaluate sequence-based models and graph neural networks (GNNs) on benchmark tasks. Furthermore, by concurrently performing eight glycan taxonomy prediction tasks, we introduce the GlycanML-MTL testbed for multi-task learning (MTL) algorithms. Experimental results show the superiority of modeling glycans with multi-relational GNNs, and suitable MTL methods can further boost model performance. We provide all datasets and source codes at https://github.com/GlycanML/GlycanML and maintain a leaderboard at https://GlycanML.github.io/project.
@@ 1 Introduction
The GlycanML benchmark accommodates two glycan representation structures, i.e., glycan tokenized sequences and glycan planar graphs. For each structure, we adopt suitable machine learning models for representation learning, where sequence encoders such as CNN [22], LSTM [24] and Transformer [53] are employed for glycan sequence encoding, and both homogeneous GNNs [29, 54, 59] and heterogeneous GNNs [20, 45, 52] are used to encode glycan graphs. We evaluate each model on all benchmark tasks to study its general effectiveness. The GlycanML benchmark also provides a testbed, namely GlycanML-MTL, for multi-task learning (MTL) algorithms, where an MTL method is asked to simultaneously solve eight glycan taxonomy prediction problems which are highly correlated. The performance on this testbed measures how well an MTL method can transfer the knowledge learned from different glycan taxonomies, e.g., transferring between species-level classification and genus-level classification.
Benchmark results show that the RGCN model [45], a typical heterogeneous GNN, performs best on most benchmark tasks, and a simple two-layer CNN can surprisingly achieve competitive performance by using only condensed sequential information of glycan structures. The MTL methods with elaborate task-reweighting strategies can further enhance the performance of glycan taxonomy prediction, showing the potential of MTL for glycan understanding. We hope the GlycanML benchmark will spark the interest of studying glycoscience with machine learning.
@@ 3.1 Glycan Taxonomy Prediction
Task definition. In the GlycanML benchmark, we study glycan taxonomy prediction on domain, kingdom, phylum, class, order, family, genus and species levels, leading to eight individual tasks. These tasks are formulated as classification problems with 4, 11, 39, 101, 210, 415, 922 and 1,737 biological categories, respectively. We report classification accuracy for each task.
@@
Task Task type #Mono. per glycan #Sample #Train/Validation/Test Metric Taxonomy prediction of Domain Classification 6.39(3.51) 13,209 11,010/1,280/919 Accuracy (%) Taxonomy prediction of Kingdom Classification 6.39(3.51) 13,209 11,010/1,280/919 Accuracy (%) Taxonomy prediction of Phylum Classification 6.39(3.51) 13,209 11,010/1,280/919 Accuracy (%) Taxonomy prediction of Class Classification 6.39(3.51) 13,209 11,010/1,280/919 Accuracy (%) Taxonomy prediction of Order Classification 6.39(3.51) 13,209 11,010/1,280/919 Accuracy (%) Taxonomy prediction of Family Classification 6.39(3.51) 13,209 11,010/1,280/919 Accuracy (%) Taxonomy prediction of Genus Classification 6.39(3.51) 13,209 11,010/1,280/919 Accuracy (%) Taxonomy prediction of Species Classification 6.39(3.51) 13,209 11,010/1,280/919 Accuracy (%) Immunogenicity prediction Binary classification 7.30(3.78) 1,320 1,026/149/145 AUPRC Glycosylation type prediction Classification 9.04(3.96) 1,683 1,356/163/164 Accuracy (%) Protein-Glycan interaction prediction Regression 6.56(4.54) 564,647 442,396/58,887/63,364 Spearman’s $\rho$
@@ 3.3 Glycosylation Type Prediction
Task definition. Given a glycan, we aim at predicting whether it forms N-glycosylation, O-glycosylation or maintains a free state, formulated as a three-way classification problem. The classification accuracy is used for evaluation.
@@ 3.4 Protein-Glycan Interaction Prediction
Benchmark dataset. This benchmark dataset is built upon 564,647 protein-glycan interactions deposited in the LectinOracle database [38]. It is desired to have a model that can well generalize to new proteins against training ones, considering the continuous discovery of new proteins by sequencing techniques. Therefore, we split the dataset based on protein sequence similarity. Specifically, we first cluster all protein sequences using MMseqs2 [47] (minimum sequence identity within each cluster: 0.5), and then we derive training, validation and test proteins by splitting all clusters with an 8:1:1 ratio. Finally, samples of protein-glycan pairs are split according to the protein splits.
@@ 4.1 Representations
In the GlycanML benchmark, we adopt two glycan representation structures, i.e., glycan tokenized sequence and glycan planar graph, as illustrated in Figure 2.
@@
Model Input Layer Hidden Layers Output Layer #Params. Sequence encoders Shallow CNN [46] 128-dim. token embedding 2 $\times$ 1D conv. layers (hidden dim.: 128; max pooling over all tokens 191.7K kernel size: 5; stride: 1; padding: 2) ResNet [22] 512-dim. token embedding 8 $\times$ residual blocks (hidden dim.: 512; attentive weighted sum over all tokens 11.4M + 512-dim. positional embedding kernel size: 3; stride: 1; padding: 1) LSTM [24] 640-dim. token embedding 3 $\times$ bidirectional LSTM layers weighted sum over all tokens 26.7M (hidden dim.: 640) + linear (output dim.: 640) + Tanh Transformer [53] 512-dim. token embedding 4 $\times$ Transformer blocks (hidden dim.: 512; linear (output dim.: 512) + Tanh 21.4M + 512-dim. positional embedding #attn. heads: 8; activation: GELU) upon [CLS] token Homogeneous GNNs GCN [29] 128-dim. node embedding 3 $\times$ GCN layers concat. mean & max pooling 67.8K GAT [54] 128-dim. node embedding 3 $\times$ GAT layers (#attn. heads: 2) concat. mean & max pooling 69.4K GIN [59] 128-dim. node embedding 3 $\times$ GIN layers concat. mean & max pooling 117.4K Heterogeneous GNNs MPNN [20] 128-dim. node & edge embedding 3 $\times$ MPNN layers Set2Set pooling [56] (#steps: 3) 4.0M RGCN [45] 128-dim. node embedding 3 $\times$ RGCN layers concat. mean & max pooling 4.2M CompGCN [52] 128-dim. node embedding 3 $\times$ CompGCN layers concat. mean & max pooling 150.4K
@@ 4.2 Baselines
We include three types of models in our benchmark, i.e., sequence encoders for modeling glycan tokenized sequences, and homogeneous and heterogeneous GNNs for modeling glycan planar graphs, with 10 baseline models in total. Their detailed architectures are provided in Table 2.
@@ 5.1 Experimental Setups
Model Domain Kingdom Phylum Class Order Family Genus Species Immuno Glycos Interaction Mean Rank Sequence encoders Shallow CNN [46] 93.76(0.66) 91.66(2.29) 86.87(0.71) 70.62(0.76) 47.01(1.86) 45.27(1.52) 38.88(1.24) 33.70(1.12) 0.776(0.267) 97.22(0.44) 0.261(0.008) 3.5 ResNet [22] 93.29(1.46) 89.52(0.72) 81.32(3.16) 65.29(1.54) 41.49(1.62) 37.36(0.99) 32.93(2.88) 26.59(1.89) 0.754(0.124) 98.55(0.58) 0.273(0.004) 6.0 LSTM [24] 92.78(1.16) 88.43(1.55) 81.61(1.36) 63.47(1.95) 41.20(2.84) 38.74(1.47) 28.87(2.77) 26.04(1.73) 0.862(0.016) 96.28(1.42) 0.280(0.001) 6.3 Transformer [53] 91.98(0.27) 87.34(1.16) 80.49(1.91) 62.13(1.47) 38.30(1.10) 33.30(0.86) 27.97(1.61) 27.49(1.20) 0.729(0.069) 95.90(1.45) 0.244(0.009) 8.5 Homogeneous GNNs GCN [29] 94.38(0.51) 92.06(0.11) 80.30(0.44) 63.26(0.64) 38.70(1.06) 34.93(0.76) 33.37(0.49) 31.01(0.87) 0.688(0.023) 95.90(0.11) 0.233(0.009) 7.2 GAT [54] 94.27(0.41) 92.56(0.25) 80.81(0.60) 62.57(2.29) 40.77(2.16) 37.50(0.91) 36.38(1.10) 34.13(0.99) 0.685(0.053) 94.63(0.39) 0.229(0.002) 6.6 GIN [59] 94.41(0.17) 92.35(0.17) 84.04(1.07) 67.54(0.31) 35.40(1.92) 40.04(0.86) 34.28(0.44) 31.85(2.19) 0.716(0.051) 97.16(0.19) 0.249(0.004) 5.1 Heterogeneous GNNs MPNN [20] 93.83(0.17) 90.39(0.94) 82.48(0.94) 66.52(1.76) 45.85(0.87) 41.60(0.64) 37.69(2.02) 33.80(1.87) 0.674(0.119) 97.41(0.90) 0.217(0.002) 5.6 RGCN [45] 94.78(0.11) 91.80(0.31) 86.94(0.78) 74.17(0.44) 47.91(2.07) 46.35(0.50) 40.30(1.09) 38.12(1.15) 0.780(0.006) 95.14(2.21) 0.262(0.005) 2.5 CompGCN [52] 93.94(0.41) 93.22(0.38) 86.40(0.76) 69.28(0.99) 45.77(0.85) 44.03(1.51) 40.70(1.73) 40.04(1.32) 0.692(0.006) 94.38(3.13) 0.257(0.004) 3.9
@@ 5.2 Benchmark Results on Single-Task Learning
In Table 3, we report the single-task performance of 10 representative glycan sequence and graph encoders on benchmark tasks, and the mean rank of each model over the whole benchmark is also presented. Based on these results, we highlight the following findings:
@@
  • Multi-relational GNNs show superiority in glycan modeling. Two typical multi-relational GNNs, i.e., RGCN and CompGCN, respectively rank first and third place in terms of mean rank. Especially, RGCN achieves the best performance on 5 out of 11 benchmark tasks. Therefore, it is beneficial to model a glycan as a multi-relational graph, where different types of glycosidic bonds are deemed as different relations between monosaccharides.
@@
  • A simple shallow CNN is surprisingly effective. It is observed that the 2-layer shallow CNN ranks second place in terms of mean rank, and it gains the second-best performance on 4 out of 11 tasks and the third-best performance on 3 out of 11 tasks. We thereby demonstrate that such a shallow CNN model is sufficient to produce informative glycan representations and lead to competitive performance, which aligns with previous findings that shallow CNNs can well model biological sequences like protein sequences [46, 60].
@@
  • It is important to utilize glycosidic bond information. We can observe clear performance gains of heterogeneous GNNs over homogeneous GNNs on glycan modeling, where in terms of mean rank, three heterogeneous GNNs rank 1st, 3rd and 5th places, while three homogeneous GNNs rank 4th, 8th and 9th places. Compared to homogeneous GNNs that regard all glycosidic bonds as the same, heterogeneous GNNs fully utilize glycosidic bond information by individually treating each type of bonds, leading to obvious benefits.
@@
Table 4: Benchmark results on multi-task learning. We report mean (std) for each experiment. Two color scales of blue denote the first and second best performance. Abbr., Acc: Accuracy.
@@
Method Domain Kingdom Phylum Class Order Family Genus Species Mean Acc (%) Backbone encoder: Shallow CNN Single-Task 93.76(0.66) 91.66(2.29) 86.87(0.71) 70.62(0.76) 47.01(1.86) 45.27(1.52) 38.88(1.24) 33.70(1.12) 63.47(0.42) N-MTL 93.25(0.38) 91.51(0.54) 84.77(0.47) 70.58(0.44) 47.62(0.41) 44.11(1.63) 39.32(1.16) 36.74(0.35) 63.49(0.55) GN [14] 92.96(0.49) 91.08(0.89) 84.95(0.69) 70.29(0.66) 48.13(1.21) 42.40(1.49) 38.74(0.93) 36.49(1.63) 63.13(0.25) TS [28] 93.58(0.50) 91.55(1.03) 85.24(0.62) 71.49(0.58) 47.08(0.60) 45.19(2.50) 40.19(1.54) 35.84(1.20) 63.77(0.19) UW [28] 92.93(0.94) 91.59(1.34) 85.56(0.94) 69.57(2.24) 45.96(2.16) 44.40(0.00) 38.19(0.22) 34.78(1.11) 62.87(0.87) DWA [37] 93.25(0.82) 92.31(0.44) 84.73(0.77) 69.06(3.36) 45.85(3.02) 41.57(2.17) 35.26(3.79) 33.04(0.95) 61.88(0.90) DTP [21] 93.54(0.41) 90.79(1.26) 83.71(0.06) 69.02(2.47) 46.50(1.24) 42.76(2.02) 40.26(1.10) 36.02(1.23) 62.83(0.45) Backbone encoder: RGCN Single-Task 94.78(0.11) 91.80(0.31) 86.94(0.78) 74.17(0.44) 47.91(2.07) 46.35(0.50) 40.30(1.09) 38.12(1.15) 65.05(0.21) N-MTL 93.04(1.80) 91.59(1.67) 85.75(2.57) 72.33(1.57) 47.99(1.75) 45.67(1.64) 40.66(1.36) 38.59(1.44) 64.45(0.40) GN [14] 93.73(0.72) 91.88(0.80) 85.38(0.66) 70.80(0.33) 46.32(1.09) 44.11(0.13) 39.79(1.36) 38.67(1.52) 63.83(0.65) TS [28] 94.52(0.23) 91.62(0.78) 86.62(0.86) 73.88(1.69) 50.63(0.33) 47.99(0.22) 44.87(0.72) 43.31(0.29) 66.68(0.23) UW [28] 93.43(0.56) 91.77(0.23) 85.60(0.41) 73.45(1.61) 48.28(1.42) 44.90(0.88) 42.15(0.71) 39.61(0.89) 64.90(0.39) DWA [37] 94.09(0.50) 91.80(1.11) 86.43(1.27) 69.42(1.25) 45.77(1.03) 44.11(1.24) 40.08(0.93) 37.47(0.44) 63.65(0.25) DTP [21] 94.02(0.66) 92.53(0.44) 85.09(0.86) 69.60(1.75) 46.17(0.70) 44.29(0.47) 41.49(1.28) 39.97(2.11) 64.15(0.25)

### v2

@@ Abstract
Glycans are basic biomolecules and perform essential functions within living organisms. The rapid increase of functional glycan data provides a good opportunity for machine learning solutions to glycan understanding. However, there still lacks a standard machine learning benchmark for glycan function prediction. In this work, we fill this blank by building a comprehensive benchmark for Glycan Machine Learning (GlycanML). The GlycanML benchmark consists of diverse types of tasks including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. Glycans can be represented by both sequences and graphs in GlycanML, which enables us to extensively evaluate sequence-based models and graph neural networks (GNNs) on benchmark tasks. Furthermore, by concurrently performing eight glycan taxonomy prediction tasks, we introduce the GlycanML-MTL testbed for multi-task learning (MTL) algorithms. Also, we evaluate how taxonomy prediction can boost other three function prediction tasks by MTL. Experimental results show the superiority of modeling glycans with multi-relational GNNs, and suitable MTL methods can further boost model performance. We provide all datasets and source codes at https://github.com/GlycanML/GlycanML and maintain a leaderboard at https://GlycanML.github.io/project.
@@ 1 Introduction
The GlycanML benchmark accommodates two glycan representation structures, i.e., glycan tokenized sequences and glycan planar graphs. For each structure, we adopt suitable machine learning models for representation learning, where sequence encoders such as CNN [24], LSTM [26] and Transformer [56] are employed for glycan sequence encoding, and both homogeneous GNNs [31, 57, 62] and heterogeneous GNNs [22, 48, 55] are used to encode glycan graphs. In addition, two state-of-the-art small molecule encoders [66, 45] are also studied for all-atom glycan modeling. We evaluate each model on all benchmark tasks to study its general effectiveness. The GlycanML benchmark also provides a testbed, namely GlycanML-MTL, for multi-task learning (MTL) algorithms, where an MTL method is asked to simultaneously solve eight glycan taxonomy prediction problems which are highly correlated. The performance on this testbed measures how well an MTL method can transfer the knowledge learned from different glycan taxonomies, e.g., transferring between species-level classification and genus-level classification. Taking a step further, we study the effect of MTL on boosting three function prediction tasks (i.e., immunogenicity, glycosylation and interaction prediction) by concurrently performing taxonomy prediction.
Benchmark results show that the RGCN model [48], a typical heterogeneous GNN, performs best on most benchmark tasks, and a simple two-layer CNN can surprisingly achieve competitive performance by using only condensed sequential information of glycan structures. The MTL methods with elaborate task-reweighting strategies can further enhance the performance of glycan taxonomy prediction, and learning the immunogenicity and glycosylation type of glycans jointly with their taxonomies also achieves benefits, showing the potential of MTL for glycan understanding. We hope the GlycanML benchmark will spark the interest of studying glycoscience with machine learning.
@@ 3.1 Glycan Taxonomy Prediction
Task definition. In the GlycanML benchmark, we study glycan taxonomy prediction on domain, kingdom, phylum, class, order, family, genus and species levels, leading to eight individual tasks. These tasks are formulated as classification problems with 4, 11, 39, 101, 210, 415, 922 and 1,737 biological categories, respectively. Taking the class imbalance into consideration, we report Macro-F1 score for each task.
@@
Task Task type #Mono. per glycan #Sample #Train/Validation/Test Metric Taxonomy prediction of Domain Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Kingdom Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Phylum Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Class Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Order Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Family Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Genus Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Species Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Immunogenicity prediction Binary classification 7.30(3.78) 1,320 1,026/149/145 AUPRC Glycosylation type prediction Classification 9.04(3.96) 1,683 1,356/163/164 Macro-F1 Protein-Glycan interaction prediction Regression 6.56(4.54) 564,647 442,396/58,887/63,364 Spearman’s $\rho$
@@ 3.3 Glycosylation Type Prediction
Task definition. Given a glycan, we aim at predicting whether it forms N-glycosylation, O-glycosylation or maintains a free state, formulated as a three-way classification problem. The Macro-F1 score is used for evaluation.
@@ 3.4 Protein-Glycan Interaction Prediction
Benchmark dataset. This benchmark dataset is built upon 564,647 protein-glycan interactions deposited in the LectinOracle database [40]. It is desired to have a model that can well generalize to new proteins against training ones, considering the continuous discovery of new proteins by sequencing techniques. Therefore, we split the dataset based on protein sequence similarity. Specifically, we first cluster all protein sequences using MMseqs2 [50] (minimum sequence identity within each cluster: 0.5), and then we derive training, validation and test proteins by splitting all clusters with an 8:1:1 ratio. Finally, samples of protein-glycan pairs are split according to the protein splits, deriving the benchmark dataset for protein-glycan interaction prediction. This splitting scheme is consistent with practical vaccine design applications where newly discovered lectins (proteins) could be used as the target for glycan binding.
@@ 4.1 Representations
In the GlycanML benchmark, we adopt two glycan-specific representation structures, i.e., glycan tokenized sequence and glycan planar graph, as illustrated in Figure 2.
@@
Model Input Layer Hidden Layers Output Layer #Params. Sequence encoders Shallow CNN [49] 128-dim. token embedding 2 $\times$ 1D conv. layers (hidden dim.: 128; max pooling over all tokens 191.7K kernel size: 5; stride: 1; padding: 2) ResNet [24] 512-dim. token embedding 8 $\times$ residual blocks (hidden dim.: 512; attentive weighted sum over all tokens 11.4M + 512-dim. positional embedding kernel size: 3; stride: 1; padding: 1) LSTM [26] 640-dim. token embedding 3 $\times$ bidirectional LSTM layers weighted sum over all tokens 26.7M (hidden dim.: 640) + linear (output dim.: 640) + Tanh Transformer [56] 512-dim. token embedding 4 $\times$ Transformer blocks (hidden dim.: 512; linear (output dim.: 512) + Tanh 21.4M + 512-dim. positional embedding #attn. heads: 8; activation: GELU) upon [CLS] token Homogeneous GNNs GCN [31] 128-dim. node embedding 3 $\times$ GCN layers concat. mean & max pooling 67.8K GAT [57] 128-dim. node embedding 3 $\times$ GAT layers (#attn. heads: 2) concat. mean & max pooling 69.4K GIN [62] 128-dim. node embedding 3 $\times$ GIN layers concat. mean & max pooling 117.4K Heterogeneous GNNs MPNN [22] 128-dim. node & edge embedding 3 $\times$ MPNN layers Set2Set pooling [59] (#steps: 3) 4.0M RGCN [48] 128-dim. node embedding 3 $\times$ RGCN layers concat. mean & max pooling 4.2M CompGCN [55] 128-dim. node embedding 3 $\times$ CompGCN layers concat. mean & max pooling 150.4K All-Atom Molecular Encoders Graphormer [66] 768-dim. node embedding 8 $\times$ Graphormer layers concat. mean & max pooling 28.5M GraphGPS [45] 512-dim. node & edge embedding 6 $\times$ GraphGPS layers concat. mean & max pooling 17.4M
@@ 4.2 Baselines
We include four types of models in our benchmark, i.e., sequence encoders for modeling glycan tokenized sequences, homogeneous and heterogeneous GNNs for modeling glycan planar graphs, and all-atom molecular encoders for modeling the all-atom molecular graphs of glycans, with 12 baseline models in total. Their detailed architectures are provided in Table 2.
@@
All-atom molecular encoders. Essentially, glycans are a kind of macromolecules with hundreds of atoms, and therefore we can directly employ all-atom molecular encoders designed for small molecules to model glycans. Among existing small molecule encoders, Graphormer [66] and GraphGPS [45] achieve prominent performance, and we include them in the proposed benchmark. We follow the default architecture of these two models in their original work, as shown in Table 2.
@@ 5.1 Experimental Setups
Model Taxonomy Immuno (AUPRC) Glycos (Macro-F1) Interaction (Spearman’s $\rho$) Weighted Mean Rank Domain (Macro-F1) Kingdom (Macro-F1) Phylum (Macro-F1) Class (Macro-F1) Order (Macro-F1) Family (Macro-F1) Genus (Macro-F1) Species (Macro-F1) Sequence Encoders Shallow CNN 0.629(0.005) 0.559(0.024) 0.388(0.024) 0.342(0.020) 0.238(0.016) 0.200(0.014) 0.149(0.009) 0.115(0.008) 0.776(0.027) 0.898(0.009) 0.261(0.008) 4.97 ResNet 0.635(0.009) 0.505(0.025) 0.331(0.061) 0.301(0.010) 0.183(0.082) 0.165(0.019) 0.112(0.018) 0.073(0.007) 0.754(0.124) 0.919(0.004) 0.273(0.004) 4.63 LSTM 0.621(0.012) 0.566(0.076) 0.413(0.036) 0.272(0.029) 0.174(0.023) 0.145(0.012) 0.098(0.016) 0.078(0.008) 0.912(0.068) 0.862(0.016) 0.280(0.001) 4.71 Transformer 0.612(0.009) 0.546(0.079) 0.316(0.014) 0.235(0.022) 0.147(0.007) 0.114(0.039) 0.065(0.001) 0.047(0.008) 0.856(0.012) 0.729(0.069) 0.244(0.009) 8.06 Homogeneous GNNs GCN 0.635(0.001) 0.527(0.006) 0.325(0.024) 0.237(0.009) 0.147(0.005) 0.112(0.010) 0.095(0.009) 0.080(0.006) 0.688(0.023) 0.914(0.011) 0.233(0.009) 7.78 GAT 0.636(0.003) 0.523(0.007) 0.301(0.014) 0.265(0.012) 0.190(0.009) 0.130(0.005) 0.125(0.010) 0.103(0.009) 0.685(0.053) 0.934(0.038) 0.229(0.002) 7.22 GIN 0.632(0.004) 0.525(0.007) 0.322(0.046) 0.300(0.027) 0.179(0.002) 0.152(0.005) 0.116(0.022) 0.105(0.011) 0.716(0.051) 0.924(0.013) 0.249(0.004) 5.59 Heterogeneous GNNs MPNN 0.632(0.007) 0.638(0.050) 0.372(0.019) 0.326(0.015) 0.235(0.046) 0.161(0.004) 0.136(0.008) 0.104(0.009) 0.674(0.119) 0.910(0.006) 0.217(0.002) 8.09 RGCN 0.633(0.001) 0.647(0.054) 0.462(0.033) 0.373(0.036) 0.251(0.012) 0.203(0.008) 0.164(0.003) 0.146(0.004) 0.780(0.006) 0.948(0.004) 0.262(0.005) 2.19 CompGCN 0.629(0.004) 0.568(0.047) 0.410(0.013) 0.381(0.024) 0.226(0.011) 0.193(0.012) 0.166(0.009) 0.138(0.014) 0.692(0.006) 0.945(0.002) 0.257(0.004) 4.28 All-Atom Molecular Encoders Graphormer 0.640(0.006) 0.468(0.054) 0.249(0.041) 0.201(0.013) 0.142(0.019) 0.112(0.009) 0.077(0.006) 0.054(0.044) 0.637 (0.062) 0.856(0.009) 0.211(0.027) 11.09 GraphGPS 0.477(0.002) 0.511(0.040) 0.314(0.022) 0.261(0.051) 0.153(0.018) 0.134(0.008) 0.105(0.006) 0.065(0.017) 0.637 (0.075) 0.883(0.032) 0.247(0.016) 9.38
@@ 5.2 Benchmark Results on Single-Task Learning
In Table 3, we report the single-task performance of 12 representative glycan encoders on benchmark tasks. We measure the comprehensive performance of a model with its weighted mean rank over all tasks, where each taxonomy prediction task owns the weight of $1/8$ and each of the other three tasks owns the weight of 1, so as to balance between different types of tasks. Based on these results, we highlight the following findings:
@@
  • Multi-relational GNNs show superiority in glycan modeling. Two typical multi-relational GNNs, i.e., RGCN and CompGCN, respectively rank first and second place in terms of weighted mean rank. Especially, RGCN achieves the best performance on 6 out of 11 benchmark tasks. Therefore, it is beneficial to model a glycan as a multi-relational graph, where different types of glycosidic bonds are deemed as different relations between monosaccharides.
@@
  • A simple shallow CNN is surprisingly effective. It is observed that the 2-layer shallow CNN ranks among the top three on 5 out of 11 tasks, and it ranks fifth place in terms of weighted mean rank. We thereby demonstrate that such a shallow CNN model is sufficient to produce informative glycan representations and lead to competitive performance, which aligns with previous findings that shallow CNNs can well model biological sequences like protein sequences [49, 63].
@@
  • It is important to utilize glycosidic bond information. We can observe clear performance gains of heterogeneous GNNs over homogeneous GNNs on glycan modeling, where in terms of weighted mean rank, three heterogeneous GNNs rank 1st, 2nd and 10th places, while three homogeneous GNNs rank 6th, 7th and 8th places. Compared to homogeneous GNNs that regard all glycosidic bonds as the same, heterogeneous GNNs fully utilize glycosidic bond information by individually treating each type of bonds, leading to obvious benefits.
  • Small molecule encoders can hardly handle glycan modeling. Though performing well on small molecule modeling, Graphormer and GraphGPS perform poorly on the GlycanML benchmark, where they rank last two places in terms of weighted mean rank. These results illustrate that the small molecule encoders originally designed to model tens of atoms cannot well model a macromolecule like a glycan with hundreds of atoms, calling for specifically designed models for all-atom glycan modeling.
@@
Table 4: Benchmark results on multi-task learning. We report mean (std) for each experiment. Two color scales of blue denote the first and second best performance.
@@
Method Domain Kingdom Phylum Class Order Family Genus Species Mean Macro-F1 Backbone encoder: Shallow CNN Single-Task 0.629(0.005) 0.559(0.024) 0.388(0.024) 0.342(0.020) 0.238(0.016) 0.200(0.014) 0.149(0.009) 0.115(0.008) 0.327(0.003) N-MTL 0.619(0.009) 0.549(0.040) 0.333(0.029) 0.344(0.015) 0.213(0.015) 0.167(0.014) 0.150(0.008) 0.121(0.010) 0.312(0.010) GN 0.646(0.043) 0.497(0.045) 0.344(0.064) 0.312(0.002) 0.233(0.042) 0.206(0.027) 0.175(0.038) 0.137(0.038) 0.319(0.012) TS 0.642(0.031) 0.609(0.059) 0.393(0.061) 0.355(0.013) 0.249(0.019) 0.187(0.009) 0.141(0.010) 0.113(0.009) 0.336(0.014) UW 0.632(0.005) 0.620(0.046) 0.409(0.025) 0.320(0.011) 0.230(0.012) 0.179(0.012) 0.152(0.001) 0.118(0.008) 0.332(0.005) DWA 0.619(0.012) 0.536(0.018) 0.354(0.013) 0.330(0.041) 0.218(0.011) 0.152(0.018) 0.138(0.015) 0.099(0.012) 0.305(0.013) DTP 0.633(0.004) 0.556(0.023) 0.341(0.041) 0.303(0.028) 0.213(0.034) 0.171(0.010) 0.145(0.011) 0.107(0.006) 0.309(0.017) Backbone encoder: RGCN Single-Task 0.633(0.001) 0.647(0.054) 0.462(0.033) 0.373(0.036) 0.251(0.012) 0.203(0.008) 0.164(0.003) 0.146(0.004) 0.360(0.009) N-MTL 0.641(0.001) 0.646(0.066) 0.402(0.042) 0.376(0.054) 0.231(0.025) 0.190(0.007) 0.154(0.018) 0.119(0.007) 0.345(0.020) GN 0.626(0.002) 0.612(0.050) 0.379(0.011) 0.364(0.017) 0.233(0.010) 0.169(0.006) 0.148(0.009) 0.119(0.006) 0.331(0.005) TS 0.640(0.003) 0.652(0.050) 0.461(0.013) 0.388(0.031) 0.267(0.024) 0.209(0.010) 0.174(0.002) 0.151(0.002) 0.368(0.016) UW 0.362(0.014) 0.568(0.045) 0.395(0.042) 0.386(0.045) 0.245(0.015) 0.199(0.009) 0.165(0.005) 0.133(0.012) 0.341(0.019) DWA 0.587(0.095) 0.616(0.112) 0.400(0.035) 0.379(0.024) 0.230(0.017) 0.177(0.023) 0.151(0.015) 0.110(0.005) 0.331(0.032) DTP 0.647(0.019) 0.637(0.083) 0.399(0.069) 0.370(0.023) 0.266(0.031) 0.228(0.039) 0.190(0.043) 0.158(0.035) 0.362(0.007)